# Preserving Privacy in Large Language Models: A Survey on Current Threats and Solutions

### v1

@@ 5.4 Machine Unlearning
Recent regulations like GDPR guarantees the "right to be forgotten", which allows individuals to withdraw their consent and request the deletion of their personal data (Zhang et al., 2024). Unfortunately, differential privacy (DP) cannot guarantee this because, for a data sample to be truly "forgotten", it must have zero influence on the training process. This requirement translates to a privacy budget of $\epsilon=0$, which would prevent the model from learning anything.
@@
Machine Unlearning (Cao & Yang, 2015) is an orthogonal field to Differential Privacy, focusing on how to eliminate the influence of specific data samples. The aim is to make the model equivalent to one trained on a dataset that excluded the unwanted entry, as illustrated in Figure 9. This is a particularly challenging issue due to our limited understanding of how individual data points affect model parameters. The key works related to this topic are summarized in Table 7 in the Appendix.
Some research attempts to trace test-time predictions back to the training data using influence functions (Koh & Liang, 2017), but these methods are costly and do not clearly reveal a single data point’s impact on model weights. In a recent study Meng et al. (2023) introduces ROME (Rank-One Model Editing), allowing factual predictions to be traced to specific layers and neurons for individual manipulation. While focused on factual associations, ROME shows potential for tracing other data types. However, Patil et al. (2023) demonstrate that deleted data can be recovered 38% of the time by examining intermediate hidden states.
Inspired by ROME, Wu et al. (2023b) introduce DEPN (Detecting and Editing Privacy Neurons). This method assume that private information, like factual information, might be encoded in specific neurons. They propose detecting these neurons and setting their weights to zero to block information flow. Results show good validation perplexity, sometimes surpassing models trained with Differential Privacy but with higher privacy risks. However, DEPN significantly reduces training time, offering a better balance between computational efficiency and performance.
An alternative approach called SISA (Sharded, Isolated, Sliced, Aggregated) proposed by Bourtoule et al. (2019), divides training data into shards, further split into slices, and trains several model versions on each shard incrementally. Parameters are saved at each step, and the models are aggregated. To unlearn data, only the model with the relevant shards needs retraining. However, this still requires some retraining.
Jang et al. (2022) propose a method focused on unlearning specific information, particularly for language models. This method negates the training loss function to maximize loss on target sequences, enabling unlearning with minimal performance loss. Sequential unlearning proves more effective than batch unlearning. Despite its effectiveness, Eldan & Russinovich (2023) show limitations in certain scenarios, such as making Llama2-7B forget about Harry Potter. The challenge lies in the high likelihood of several candidate tokens related to Harry Potter, necessitating many gradient descent steps. To address this, they map Harry Potter terms to generic equivalents and then compare predictions from a baseline model with mapped input and a reinforced model trained on Harry Potter with the original input. By reducing the likelihood of tokens predicted by the reinforced model, they fine-tune the baseline model using generic labels. This results in the model losing familiarity with the target knowledge, with minimal leakages resembling secondary knowledge. Performance drops vary by task but remain generally tolerable.
@@
Unlearning can also be achieved with Reinforcement Learning through an approach called DeMemorization (Kassem et al., 2023). This method takes samples from the pre-training dataset, divides them into prefix and suffix pairs, inputs the prefix, and uses a negative similarity metric to evaluate how different the generated suffix is from the real one. This metric serves as a reward signal, encouraging the language model to develop a paraphrasing policy that generates different tokens to minimize memorization. DeMemorization achieves better performance on downstream tasks because it doesn’t remove samples from the dataset. However, for the same reason, it shows less robust privacy performance compared to actual unlearning methods, though the privacy performance remains high.
@@
For large language models, In-Context Learning can be used for unlearning, as demonstrated by ICUL (In-Context UnLearning) (Pawelczyk et al., 2023). This technique builds a context where the data point to be unlearned is included but with a flipped label (e.g., changing a positive label to negative) alongside correctly labeled examples. Compared to Jang et al. (2022), this approach achieves similar performance for both unlearning and the primary task without requiring any parameter updates. While this method was tested on a binary sentiment classification task, theoretically, it could be applied to other tasks by incorrectly labeling categorical samples, though no experiments have been conducted in such contexts to our knowledge.
@@ 6 Tools & Frameworks
After discussing the different approaches one may take to tackle the problem of privacy-preserving LLMs, we briefly see what frameworks and libraries are already in place for researchers and practitioners who are interested in the topic.
@@
Differential Privacy
We have seen that Differential Privacy has been considered a gold standard for privacy protection in machine learning for years now, with several techniques to enforce it and big players using it in very used applications. Naturally, this means that there is a high number of frameworks to implement it, some of which are open source. Most of these frameworks, like most of the research in the field, focus on differentially private training. The most important ones are:
@@
  • Google DP: Provides a broad set of DP tools in Java, Go and C++. (1) Privacy on Beam: an end-to-end DP framework built on top of Apache Beam. (2) Three "DP building block" libraries in C++, Go, and Java implement basic noise addition primitives and DP aggregations. Privacy on Beam is implemented using these libraries. (3) A stochastic tester is used to help catch regressions that could make the DP no longer hold. (4) A DP accounting library, used for tracking privacy budget. (5) A command line interface for running differentially private SQL queries with ZetaSQL. DP Auditorium is a library for auditing differential privacy guarantees.
@@
Transformers
The main focus of this work is LLMs, which are based on the transformers architecture (Vaswani et al., 2023). Most transformers nowadays are implemented with the Transformers Library by HuggingFace. This library quickly became a cornerstone in the field of AI because it provides most state-of-the-art models already pre-trained and ready to be used in a plug-and-play fashion, allowing researchers to fast and efficient prototyping. It provides tools for pre-training and fine-tuning like SFT for supervised learning and TRL for reinforcement learning. The main reason for the success of this library is its versatility: there are also integrations for distributed training with Accelerate and DeepSpeed and it even supports Parameter-Efficient Fine-Tuning with an ad-hoc library called PEFT, which are all fundamental elements when you use big models like modern LLMs. There has been some work to try and integrate DP libraries and Transformers, and in particular there are two public implementations of Opacus with Transformers: private-transformers and dp-transformers which both provide implementations of some recent LLMs and other models with the integration of Opacus for Differential Privacy.
Federated Learning
Federated learning is standard when the task requires handling sensitive data, as it makes it possible to keep the data on-site without sending them to a central server and therefore gives a (false §3.3) sense of security. There are several frameworks for federated learning. Let’s quickly see the most famous ones:
@@
  • pfl: Apple’s pfl library is a federated learning library specifically devised for quick research simulations. It also provides tight integration with both local and central DP mechanisms, allowing quick prototyping of differentially private federated networks.
@@
7 Discussion and Conclusion

### v2

@@ 5.4 Machine Unlearning
Recent regulations like GDPR guarantees the "right to be forgotten", which allows individuals to withdraw their consent and request the deletion of their personal data (Zhang et al., 2024b). Unfortunately, differential privacy (DP) cannot guarantee this because, for a data sample to be truly "forgotten", it must have zero influence on the training process. This requirement translates to a privacy budget of $\epsilon=0$, which would prevent the model from learning anything.
@@
Machine Unlearning (Cao & Yang, 2015) is an orthogonal field to Differential Privacy, focusing on how to eliminate the influence of specific data samples. The key works related to this topic are summarized in Table 7 in the Appendix. The aim is to make the model equivalent to one trained on a dataset that excluded unwanted entries, as illustrated in Figure 9. While test-time prediction to measure the influence of training data have been proposed defining influence functions (Koh & Liang, 2017), this technique does not reveal the impact of individual data point model parameters. Although training strategies designed to ensure unlearning have been proposed and tested for different architectures (Bourtoule et al., 2019), most of the unlearning solutions for LLMs adopt the idea of modifying model parameters after the training phase.
Jang et al. (2022) propose a method focused on unlearning specific information, particularly for language models. This method negates the training loss function – performing gradient ascent – to maximize loss on target sequences, enabling unlearning with minimal performance loss on unrelated sequences. Sequential unlearning proves more effective than batch unlearning. Wang et al. (2024c) propose to perform the unlearning only of problematic spans in the sentence, rather than on entire instances to minimize the impact on performance. Despite its effectiveness, Eldan & Russinovich (2024) show limitations of gradient ascent in certain scenarios, such as making Llama2-7B forget about Harry Potter. The challenge lies in the high likelihood of several candidate tokens related to Harry Potter, necessitating many gradient descent steps. To address this, they map Harry Potter terms to generic equivalents and then compare predictions from a baseline model with mapped input and a reinforced model trained on Harry Potter with the original input. By reducing the likelihood of tokens predicted by the reinforced model, they fine-tune the baseline model using generic labels. This results in the model losing familiarity with the target knowledge, with minimal leakages resembling secondary knowledge. Performance drops vary by task but remain generally tolerable. Recent findings suggest that gradient ascent should be combined with other techniques to be effective while maintaining model utility: Yao et al. (2024a) describe a unified unlearning framework for LLMs and find that gradient ascent is more robust when combined with gradient descent on a retain dataset.
@@
Unlearning can also be achieved with Reinforcement Learning through an approach called DeMemorization (Kassem et al., 2023). This method takes samples from the pre-training dataset, divides them into prefix and suffix pairs, inputs the prefix, and uses a negative similarity metric to evaluate how different the generated suffix is from the real one. This metric serves as a reward signal, encouraging the language model to develop a paraphrasing policy that generates different tokens to minimize memorization. DeMemorization achieves better performance on downstream tasks because it doesn’t remove samples from the dataset. However, for the same reason, it shows less robust privacy performance than actual unlearning methods, though the privacy performance remains high.
@@
For large language models, In-Context Learning can be used for unlearning, as demonstrated by ICUL (In-Context UnLearning) (Pawelczyk et al., 2024). This technique builds a context where the data point to be unlearned is included but with a flipped label (e.g., changing a positive label to a negative) alongside correctly labelled examples. Compared to Jang et al. (2022), this approach achieves similar performance for both unlearning and the primary task without requiring any parameter updates. While this method was tested on a binary sentiment classification task, theoretically, it could be applied to other tasks by incorrectly labelling categorical samples, though no experiments have been conducted in such contexts to our knowledge.
Model Editing can also be used, in principle, to perform unlearning of private information. Building on the idea that the linear layers in the Transformer architecture can be interpreted as key-value memories that store information (Geva et al., 2021), different model editing techniques have been developed to modify – without retraining – these specific parameters. Meng et al. (2022) introduce ROME (Rank-One Model Editing), allowing factual predictions to be traced to specific layers for individual manipulation. Furthermore, a similar approach called MEMIT (Meng et al., 2023) proves to effectively modify multiple examples without affecting the utility of the model. While grey-box attacks that examine intermediate hidden states can still recover deleted data 38% of the time (Patil et al., 2024), and being mostly focused on factual associations, ROME and MEMIT show potential for tracing other data types, substituting the private information with an anonymous one.
With a more specific focus on privacy, Wu et al. (2023) introduce DEPN (Detecting and Editing Privacy Neurons) inspired by ROME. This method assume that private information, like factual information, might be encoded in specific neurons. They propose detecting these neurons and setting their weights to zero to block information flow. Results show good validation perplexity, sometimes surpassing models trained with Differential Privacy but with higher privacy risks. However, DEPN significantly reduces training time, offering a better balance between computational efficiency and performance. Venditti et al. (2024) focus on the idea of breaking the association between an individual and their private information inadvertently included in the training data. Their method is called PAE and is based on MEMIT: PAE aims to make multiple edits feasible, both in a sequential and batch fashion, by taking into account the norm of the update matrix with respect to the original weights without sacrificing the model’s utility.
5.5 Homomorphic Encryption
Homomorphic Encryption Definition: Homomorphic Encryption (HE) is a cryptographic technique that allows computations to be performed directly on encrypted data without the need to decrypt it first. The result of these computations, when decrypted, matches the result of operations performed on the plaintext data.
Homomorphic Encryption (HE) offers a theoretical framework for preserving privacy in machine learning by enabling computations on encrypted data. In the context of Large Language Models (LLMs), HE can be considered for performing inference on sensitive inputs without exposing the raw data to the model provider. This could be particularly useful when the LLM is hosted on external servers or cloud services, and users wish to keep their input data confidential. Regarding training, HE is likely to not be able to prevent memorization as the patterns in the plaintext that are usually memorized by LLMs should be preserved by the encryption step.
The application of Homomorphic Encryption to LLMs aims to address privacy concerns during the inference phase by ensuring that sensitive user inputs remain encrypted throughout the computational process. Users encrypt their queries using an HE scheme and send the encrypted data to the server hosting the LLM. The server performs computations directly on the encrypted data and returns an encrypted result, which the user can decrypt to obtain the plaintext output. While this approach theoretically provides strong privacy guarantees, several significant challenges hinder its practical implementation with LLMs.
First, computational overhead and latency are substantial. HE schemes, especially those capable of supporting the complex operations required by deep neural networks, introduce significant computational overhead. Operations on encrypted data are significantly slower than their plaintext counterparts, often by several orders of magnitude. This performance penalty makes real-time inference with LLMs impractical using current HE technologies. Second, model complexity and compatibility pose challenges. LLMs typically consist of billions of parameters and involve complex architectures with non-linear activation functions, such as GELU or Softmax, and operations like layer normalization and attention mechanisms. Many of these operations are not directly compatible with standard HE schemes, which primarily support arithmetic operations (addition and multiplication). Approximating these non-linear functions with polynomial functions compatible with HE can lead to a loss in model accuracy and additional computational complexity. Additionally, memory and storage requirements increase significantly. Encrypted data under HE schemes can be much larger than plaintext data due to encryption overhead. This ciphertext expansion results in higher memory and storage requirements on both client and server sides, making it challenging to handle large encrypted inputs and model parameters. Finally, there is limited support for training. While HE can, in theory, be applied to both training and inference, the vast computational demands make training LLMs under HE impractical with current technology. Consequently, most considerations of HE in the context of LLMs focus solely on the inference phase.
Current Research and Limitations
Research into applying Homomorphic Encryption to neural networks has shown some success with smaller models and simpler tasks. For instance, Gilad-Bachrach et al. (2016) introduced CryptoNets, demonstrating the feasibility of performing inference on encrypted data with neural networks for basic classification tasks like recognizing handwritten digits (MNIST dataset). Similarly, Al Badawi et al. (2021) proposed optimizations for running convolutional neural networks under HE but were limited to relatively small models and datasets. Extending these approaches to LLMs presents additional challenges.
Recent efforts have attempted to apply HE to Transformer models and LLMs. Hao et al. (2022) proposed Iron, a method that modifies Transformers to be more compatible with HE by approximating non-linear operations with polynomial functions, aiming to reduce computational complexity while preserving accuracy. Pang et al. (2024) introduced BOLT, which provides privacy-preserving and efficient inference for Transformers by reformulating attention computation to be more HE-friendly. Similarly, Chen et al. (2022a) presented THE-X, a framework that redesigns critical components of the Transformer architecture to suit HE constraints, optimizing encoding schemes to minimize performance degradation. Zheng et al. (2023) developed Primer, focusing on fast private Transformer inference on encrypted data by introducing optimizations like efficient polynomial approximations and streamlined encrypted computations.
Recently, Hou et al. (2023) introduced CipherGPT, a framework designed for secure two-party GPT inference that addresses privacy concerns in LLMs. CipherGPT enables a client (holding the input data) and a server (hosting the model) to collaboratively perform GPT inference without revealing the client’s input or the server’s model parameters to each other. This is achieved through the use of cryptographic protocols such as secure multiparty computation and homomorphic encryption, optimized specifically for GPT models.
CipherGPT introduces several technical innovations to improve the efficiency and practicality of secure inference with large models like GPT. These include a customized secure matrix multiplication protocol using subfield vector oblivious linear evaluation (sVOLE), an efficient method for computing the GELU activation function using spline-based approximations, and protocols for secure top-$K$ selection and sampling required for the text generation process. By addressing the computational and communication overheads typically associated with homomorphic encryption, CipherGPT demonstrates the feasibility of performing privacy-preserving inference with large-scale language models. However, despite these advancements, CipherGPT still faces challenges related to computational overhead and scalability. The secure protocols, although optimized, introduce latency and resource consumption that may hinder real-time applications, especially with larger models. Additionally, the approximations used for non-linear functions may affect the model’s performance and the quality of the generated text. Implementing CipherGPT also requires specialized expertise in cryptography and deep learning, which can be a barrier to widespread adoption.
While these works represent significant progress, they are often constrained by computational overhead and are applied to smaller models or simplified versions of Transformers. One major challenge is approximating non-linear functions. LLMs rely heavily on non-linear operations not directly supported by HE schemes. Researchers have explored using polynomial approximations of these functions, but such approximations can degrade model performance. For example, approximating the Softmax function used in attention mechanisms with a polynomial can lead to less accurate attention scores, affecting the overall quality of the generated text.
Another issue is the pronounced performance trade-off between privacy and utility. With LLMs, due to their size and complexity, this trade-off is more significant. While HE provides strong privacy by keeping data encrypted, the resulting computational costs and potential loss in accuracy make it less feasible for practical applications involving LLMs.
Furthermore, there is a lack of practical implementations for full-scale LLMs. Although recent works like Rho et al. (2024) have provided implementations for encoder-only models, they still face scalability issues when applied to large models. Li et al. (2023a) introduced MPCFormer, which enables private Transformer inference using Secure Multiparty Computation (MPC) instead of HE, optimizing the Transformer architecture to be compatible with MPC protocols. While MPCFormer shows improved performance over traditional HE methods, it still encounters challenges with communication overhead and scalability when applied to large-scale LLMs.
While Homomorphic Encryption presents an attractive theoretical solution for preserving privacy during inference with Large Language Models, current technological limitations render it impractical for real-world applications involving LLMs. The substantial computational overhead, memory requirements, and challenges in handling complex non-linear operations make HE unsuitable for deployment with current models. Alternative privacy-preserving techniques, such as Secure Multiparty Computation (SMC) and Trusted Execution Environments (TEEs), face similar challenges when applied to LLMs. Differential Privacy (as discussed in §5.2) and Federated Learning (§5.1) currently offer more practical approaches for preserving privacy in the context of LLMs, balancing efficiency with privacy guarantees. Ongoing research in cryptographic techniques and hardware acceleration may, in the future, alleviate some of the computational burdens associated with Homomorphic Encryption. Until such advancements are realized, HE remains a topic of interest in theoretical discussions but is not a practical solution for privacy preservation in Large Language Models.
@@ 6 Tools & Frameworks
After discussing the approaches one may take to tackle the problem of privacy-preserving LLMs, we briefly see what frameworks and libraries are already in place for researchers and practitioners interested in the topic.
@@
Differential Privacy.
We have seen that Differential Privacy has been considered a gold standard for privacy protection in machine learning for years, with several techniques to enforce it and big players using it in very used applications. Naturally, this means that there are many frameworks to implement, some of which are open-source. Most of these frameworks, like most of the research in the field, focus on differentially private training. Here are presented the most prominent ones, for the tabular version of this list, refer to 1:
@@
  • Google DP: Provides a broad set of DP tools in Java, Go and C++. (1) Privacy on Beam: an end-to-end DP framework built on top of Apache Beam. (2) Three "DP building block" libraries in C++, Go, and Java implement basic noise addition primitives and DP aggregations. Privacy on Beam is implemented using these libraries. (3) A stochastic tester is used to help catch regressions that could make the DP no longer hold. (4) A DP accounting library is used to track the privacy budget. (5) A command line interface for running differentially private SQL queries with ZetaSQL. DP Auditorium is a library for auditing differential privacy guarantees.
@@
Transformers.
The main focus of this work is LLMs, which are based on the transformers architecture (Vaswani et al., 2017). Most transformers nowadays are implemented with the Transformers Library by HuggingFace. This library quickly became a cornerstone in AI because it provides most state-of-the-art models already pre-trained and ready to be used in a plug-and-play fashion, allowing researchers to fast and efficient prototyping. It provides tools for pre-training and fine-tuning like SFT for supervised learning and TRL for reinforcement learning. The main reason for the success of this library is its versatility: there are also integrations for distributed training with Accelerate and DeepSpeed and it even supports Parameter-Efficient Fine-Tuning with an ad-hoc library called PEFT, which are all fundamental elements when you use big models like modern LLMs. There has been some work to try and integrate DP libraries and Transformers, and in particular there are two public implementations of Opacus with Transformers: private-transformers and dp-transformers which both provide implementations of some recent LLMs and other models with the integration of Opacus for Differential Privacy.
Federated Learning.
Federated learning is standard when the task requires handling sensitive data, as it makes it possible to keep the data on-site without sending them to a central server and, therefore, gives a (false §3.3) sense of security. There are several frameworks for federated learning. Let’s quickly see the most famous ones:
@@
  • pfl: Apple’s pfl library is a federated learning library specifically devised for quick research simulations. It also provides tight integration with local and central DP mechanisms, allowing quick prototyping of differentially private federated networks.
@@
7 Limitations and Future Directions
In this section, we provide a discussion on current limitations and future directions for preserving privacy in large language models, analyzing all the different approaches to preserving privacy proposed in this survey.
Preserving Privacy in Data.
Incorporating privacy directly within data before it is input into a large language model offers a practical alternative to more complex, model-based privacy-preserving techniques. For example, the large-scale pretraining of prominent models like LLaMA3 and OLMo leverages data anonymization techniques, applying privacy measures at the data level before it enters the LLM. However, given the vast scale of pretraining data, these projects have largely relied on basic regular expressions to filter out PII and streamline the preprocessing (Dubey et al., 2024; Soldaini et al., 2024). While efficient, this approach is quite limited as it lacks comprehensive privacy protection and does not offer formal privacy guarantees such as those provided by DP. As Subramani et al. (2023) notes, “Running an out-of-the-box tool such as Presidio is the bare minimum that should be applied to all new and existing corpora”. To support progress in this area, sections 4.1 and 6 of this survey present practical insights for researchers and practitioners aiming to build the next generation of privacy-preserving LLMs. While current models primarily rely on regex-based filtering for pretraining data, several approaches discussed in section 4.1, including metric Differential Privacy methods, can be scaled and applied to enhance privacy preservation in LLMs. We believe these methods hold significant untapped potential for advancing privacy-preserving techniques in data handling for LLMs and that we are just beginning to explore their full possibilities.
Preserving Privacy in Model.
Model-based approaches based on DP offer mathematically-proven privacy guarantees within defined privacy parameters $(\epsilon,\delta)$, but they come with significant computational costs and require substantial, often complex, modifications to models and training procedures. Of all the approaches discussed in section 5, DP is considered the gold standard for privacy-preserving machine learning, as it is the only method that provides a formal mathematical privacy guarantee. However, this assurance typically comes at the expense of some utility loss, especially when DP is applied throughout the entire learning pipeline (Tramèr et al., 2024b; Feng & Tramèr, 2025). DP’s applicability spans data preparation, training, fine-tuning, and inference. In contrast, the privacy guarantee provided by other approaches like Machine Unlearning (§5.4) is only empirical but necessary to implement the "right to be forgotten" of privacy laws like GDPR (Zhang et al., 2024b). Future research could focus on building a rigorous mathematical basis for Machine Unlearning, developing techniques to seamlessly combine it with DP, and implementing accessible interfaces that allow users to specify data they wish to be erased from the model.
Despite the extensive application scope of Differential Privacy, it assumes a structure of private data where every bit of language can potentially contain secrets that are not meant to be shared indiscriminately. Differential Privacy cannot directly preserve secrets – information that can be shared only to the correct group of people (Brown et al., 2022b). The capability of LLMs to keep secrets is still little explored, with the development of the first benchmarks to assess the ability of such models to disclose private information only to users who should be allowed to access them (Mireshghallah et al., 2024; Priyanshu et al., 2023; Wang et al., 2024a). Future research should focus on advancing DP to selectively protect specific language elements, enhancing utility by applying privacy constraints only where necessary. As LLMs are increasingly integrated as assistants in applications, functioning as agents among groups of users who wish to share certain information while keeping other details private, addressing this issue will become more critical.
Looking ahead, applying DP to the training, tuning, or conditioning of larger, more capable models presents both a challenge and an opportunity. Although most current studies focus on smaller models, the advancement of quantization and Parameter Efficient Fine-Tuning (PEFT) techniques promises to make DP viable in larger models as well. These advancements reduce computational overhead and have facilitated DP-tuning in smaller models, such as RoBERTa and GPT-2 via LoRA, which could soon extend to modern models like Llama-3 and Mistral with negligible impact on computational resources (§5.1.3). We believe that the next step in this field will be the development of privacy-preserving LLMs that leverage these techniques, paired with state-of-the-art models. With the existing building blocks (techniques and models) in place, this development hinges primarily on computational resources for private fine-tuning. Our ideal vision includes models pre-trained entirely with DP, which would significantly mitigate privacy risks from the outset. However, we recognize that fully DP-compliant pre-training remains challenging, as it demands significant computational resources and can reduce model effectiveness. A practical compromise could involve pre-training the model using data anonymization techniques, which are less computationally intensive (as discussed in Section 4.1) while reserving DP methods for the subsequent stages of instruction tuning and alignment. In the short term, we believe this approach offers a reasonable balance.
Federated Learning (§5.3) offers another promising avenue for enhancing privacy in LLMs, albeit with limitations. Although FL with Central DP is feasible, constraints on device capacity mean that edge devices, such as mobile phones, often lack the processing power to maintain a locally stored LLM. This limitation is slowly diminishing, as illustrated by devices capable of running small-scale models like Gemini-Nano locally (Team et al., 2024). With the ongoing miniaturization and optimization of LLMs, we anticipate that FL-driven, DP-compliant training on local devices will become increasingly feasible. Nevertheless, implementing a local DP model remains challenging due to the high dimensionality of the data and model size; PEFT techniques provide some relief by minimizing parameter sharing, but more research is needed to refine utility and privacy outcomes within this framework.
Inference under DP constraints has shown initial success, suggesting that LLMs may offer privacy-preserving inference without requiring task-specific fine-tuning (§5.2). While current research in DP-constrained inference is limited, early results are promising, especially as models become more compact without sacrificing capability. In practical terms, this progress allows for privacy-preserving inference in cases where offsite models are preferable or where specific fine-tuning is unnecessary. Future work could explore larger models, potentially introducing DP to inference on locally stored models, where the option of performing inference without any data sharing would be highly advantageous.
Despite the appeal of Homomorphic Encryption (HE) for privacy-preserving inference with LLMs, significant challenges limit its practicality. As outlined in §5.5, HE introduces considerable computational overhead and latency, making real-time inference with large models unfeasible. Issues with approximating non-linear functions can degrade performance, while ciphertext expansion adds memory demands that complicate deployment on limited-resource systems. Recent advancements like CipherGPT (Hou et al., 2023) have optimized HE for LLMs, yet scalability, performance trade-offs, and complexity remain concerns. Future research should aim to reduce overhead through more efficient protocols, improve non-linear approximations, and explore hybrid methods that combine HE with other privacy techniques.
In conclusion, our findings reveal that while privacy-preserving techniques for ML are advancing, efficiency remains a central focus, as existing methods often involve trade-offs in training speed and model performance. Recent innovations have brought privacy-preserving approaches closer to practical deployment in public-facing models, reducing these trade-offs and making DP increasingly accessible. We are confident that this trend will persist, paving the way for the widespread use of privacy-preserving ML as a practical alternative in the near future.
8 Conclusion